# The CFT of Sen’s Formulation of Chiral Gauge Fields

### v2

@@ Appendix A A Trinity of Energy-Momentum Tensors
This implies that $(1+\bar{\star})(\bar{\delta}\bar{\star})\star Q=0$ and hence $(\bar{\delta}\bar{\star})Q=\tfrac{1}{2}(1-\bar{\star})(\bar{\delta}\bar{\star})Q$. Thus we have
@@
Next we need to compute $\bar{\delta}{M}$. First we observe that $(1+\bar{\star}){M}=0$ and hence
@@
Thus we have $(1-\bar{\star})(\bar{\delta}\bar{\star}){M}=0$ and hence $\bar{\delta}{M}=-\tfrac{1}{2}(\bar{\delta}\bar{\star}){M}+\bar{\Xi}$ for some $\bar{\Xi}$ that is anti-self-dual: $\Xi=\tfrac{1}{2}(1-\bar{\star})\Xi$. It follows that
$=2\Upsilon\wedge M(Q)+Q\wedge\Xi(Q),$ (134)
as (anti-)self-dual forms have vanishing wedge products with themselves. Now we note that
@@
and hence
@@
$=(1-\star)\left(\frac{1}{2}(\bar{\delta}\bar{\star})Q+\bar{\Xi}(Q)+M(\Upsilon)\right)$
$=(1-\star)\left(\frac{1}{2}(\bar{\delta}\bar{\star})Q+\Xi(Q)+M(\Upsilon)\right)$ (136)
where in the second line we used the fact that M{M} vanishes on anti-self-dual forms. Since $\frac{1}{2}(\bar{\delta}\bar{\star})Q$ is anti-self dual we find that
$\Xi(Q)=-\frac{1}{2}(\bar{\delta}\bar{\star})Q-M(\Upsilon),$ (137)
which is indeed anti-self-dual. However $\Upsilon$ is not determined. Therefore we find
$\bar{\delta}\left(Q\wedge{M}(Q)\right)$ $=2\Upsilon\wedge M(Q)-\frac{1}{2}Q\wedge(\bar{\delta}\bar{\star})Q-Q\wedge M(\Upsilon)$
$=-\frac{1}{2}Q\wedge(\bar{\delta}\bar{\star})Q+\Upsilon\wedge M(Q).$ (138)
To continue we need to determine $\Upsilon$. To do this, we first look at the case $\Upsilon=0$. Similarly to the case above, all this is leads to the same energy-momentum tensor that we would get from
$S_{\Upsilon=0}$ $=-\int\frac{1}{2}dP\wedge\bar{\star}dP-\frac{1}{2}(Q+\bar{\star}Q)\wedge dP+\frac{1}{8}Q\wedge\bar{\star}Q,$ (139)
@@
where $\bar{\Theta}^{\prime}_{\mu\nu}$ is the contribution arising from $\Upsilon$ and now $X_{\mu}\bar{\cdot}Y_{\nu}=X_{\mu\lambda_{1}...\lambda_{2k}}Y_{\nu\rho_{1}...\rho_{2k}}\bar{g}^{\lambda_{1}\rho_{1}}...\bar{g}^{\lambda_{2k}\rho_{2k}}$ and $X\bar{\cdot}Y=X_{\lambda_{1}...\lambda_{2k+1}}Y_{\rho_{1}...\rho_{2k+1}}\bar{g}^{\lambda_{1}\rho_{1}}...\bar{g}^{\lambda_{2k+1}\rho_{2k+1}}$. A little calculation shows that this can be written as
$\bar{\Theta}_{\mu\nu}=-\frac{1}{(2k)!}G_{\mu}\bar{\cdot}G_{\nu}+\frac{1}{4}\frac{1}{(2k)!}(dP-\bar{\star}dP)_{\mu}\bar{\cdot}(dP-\bar{\star}dP)_{\nu}-\bar{\Theta}^{\prime}_{\mu\nu}.$ (141)
@@
To fix this we need to consider $\Upsilon\neq 0$. This gives an additional variation
$\delta_{\hat{\Upsilon}}S$ $=\int\frac{1}{2}\Upsilon\wedge(dP-\bar{\star}dP)+\frac{1}{4}\hat{\Upsilon}\wedge M(Q).$ (142)
A natural choice is $\Upsilon=\tfrac{1}{2}(\bar{\delta}\bar{\star})M(Q)$, which we have seen is self-dual. This means that
@@
$=(\bar{\delta}\bar{\star})F$ (143)
$\bar{\delta}{M}(Q)$ $=-\frac{1}{2}(\bar{\delta}\bar{\star}){M}(Q)-\frac{1}{2}(\bar{\delta}\bar{\star})Q-\frac{1}{2}M((\bar{\delta}\bar{\star})M(Q)).$ (144)
This expression for $\bar{\delta}M$ also appeared in Lambert:2019diy for the case $\bar{g}=\eta$. In this case
$\delta_{\hat{\Upsilon}}S$ $=-\int\frac{1}{4}(dP-\bar{\star}dP)\wedge(\bar{\delta}\bar{\star})M(Q)+\frac{1}{8}M(Q)\wedge(\bar{\delta}\bar{\star})M(Q).$ (145)
@@
$\bar{\Theta}^{\prime}_{\mu\nu}=-\frac{1}{4}\frac{1}{(2k)!}\left((dP-\bar{\star}dP)_{\mu}\bar{\cdot}M(Q)_{\nu}+(dP-\bar{\star}dP)_{\nu}\bar{\cdot}M(Q)_{\mu}+M_{\mu}(Q)\bar{\cdot}M_{\nu}(Q)\right).$ (146)
@@
Here we see that the first term is conserved on-shell whereas the second vanishes.
@@
The action is invariant under such diffeomorphisms, up to possible boundary terms which we will discard, we observe that

### v3

@@ Appendix A A Trinity of Energy-Momentum Tensors
This implies that $(1+\bar{\star})(\bar{\delta}\bar{\star})Q=0$ and hence $(\bar{\delta}\bar{\star})Q=\tfrac{1}{2}(1-\bar{\star})(\bar{\delta}\bar{\star})Q$. Thus we have
@@
Next we need to compute $\bar{\delta}{M}$. We observe that $(1+\bar{\star}){M}=0$ and hence
@@
Thus we have $(1-\bar{\star})(\bar{\delta}\bar{\star}){M}=0$ and hence
$\bar{\delta}{M}=-\frac{1}{2}(\bar{\delta}\bar{\star}){M}+\Xi.$ (134)
for some $\Xi$ that is anti-self-dual: $\Xi=\tfrac{1}{2}(1-\bar{\star})\Xi$. To continue, we note that
@@
which must be maintained when we vary $\bar{g}$. This leads to
@@
$=(1-\star)\left(\frac{1}{2}(\bar{\delta}\bar{\star})Q+\Upsilon-\frac{1}{2}(\bar{\delta}\bar{\star}){M}(Q)+\Xi(Q)+M(\Upsilon)\right)$
$=(1-\star)\left(-\frac{1}{2}(\bar{\delta}\bar{\star}){M}(Q)+\frac{1}{2}(\bar{\delta}\bar{\star})Q+\Xi(Q)\right),$ (136)
where in the second line we used the fact that M{M} vanishes on anti-self-dual forms and in the fourth line that $\Upsilon+M(\Upsilon)=\star(\Upsilon+M(\Upsilon))$. Thus we learn that $\frac{1}{2}(\bar{\delta}\bar{\star}){M}(Q)+\frac{1}{2}(\bar{\delta}\bar{\star})Q+\Xi(Q)$ is also self-dual with respect to $\star$ and can be written as
$-\frac{1}{2}(\bar{\delta}\bar{\star}){M}(Q)+\frac{1}{2}(\bar{\delta}\bar{\star})Q+\Xi(Q)=\Pi+M(\Pi),$ (137)
for some $\Pi$ which is self-dual with respect to $\bar{\star}$. Projecting on the self-dual and anti-self-dual parts we find
$\Pi=-\frac{1}{2}(\bar{\delta}\bar{\star}){M}(Q),\qquad M(\Pi)=\frac{1}{2}(\bar{\delta}\bar{\star})Q+\Xi(Q),$ (138)
which allows us to identify
$-\frac{1}{2}M((\bar{\delta}\bar{\star}){M}(Q))=\frac{1}{2}(\bar{\delta}\bar{\star})Q+\Xi(Q).$ (139)
From this we can determine $\Xi$:
$\Xi(Q)=-\frac{1}{2}(\bar{\delta}\bar{\star})Q-\frac{1}{2}M((\bar{\delta}\bar{\star}){M}(Q)),$ (140)
and hence $\bar{\delta}M$:
$\bar{\delta}{M}=-\frac{1}{2}(\bar{\delta}\bar{\star})Q-\frac{1}{2}(\bar{\delta}\bar{\star}){M}-\frac{1}{2}M((\bar{\delta}\bar{\star}){M}(Q)).$ (141)
A similar expression also appeared in Lambert:2019diy for the case $\bar{g}=\eta$.
Let us return to the variation of the action and note that
$=2\Upsilon\wedge M(Q)+Q\wedge\Xi$ (142)
where we have used the fact that (anti-)self-dual forms have vanishing wedge products with themselves. Thus we find
$\bar{\delta}\left(Q\wedge{M}(Q)\right)$ $=2\Upsilon\wedge M(Q)-\frac{1}{2}Q\wedge(\bar{\delta}\bar{\star})Q-\frac{1}{2}Q\wedge M((\bar{\delta}\bar{\star}){M}(Q))$
$=2\Upsilon\wedge M(Q)-\frac{1}{2}Q\wedge(\bar{\delta}\bar{\star})Q-\frac{1}{2}(\bar{\delta}\bar{\star}){M}(Q)\wedge M(Q).$ (143)
Finally we need to determine $\Upsilon$. To do this, we first look at the case $\Upsilon=0$ and drop the last term in (A). Similarly to the case above, this is leads to the same energy-momentum tensor that we would get from
$S^{\prime}$ $=-\int\frac{1}{2}dP\wedge\bar{\star}dP-\frac{1}{2}(Q+\bar{\star}Q)\wedge dP+\frac{1}{8}Q\wedge\bar{\star}Q,$ (144)
@@
where $\bar{\Theta}^{\prime}_{\mu\nu}$ is the contribution arising from $\Upsilon$ and the last term in (A). Note that now $X_{\mu}\bar{\cdot}Y_{\nu}=X_{\mu\lambda_{1}...\lambda_{2k}}Y_{\nu\rho_{1}...\rho_{2k}}\bar{g}^{\lambda_{1}\rho_{1}}...\bar{g}^{\lambda_{2k}\rho_{2k}}$ and $X\bar{\cdot}Y=X_{\lambda_{1}...\lambda_{2k+1}}Y_{\rho_{1}...\rho_{2k+1}}\bar{g}^{\lambda_{1}\rho_{1}}...\bar{g}^{\lambda_{2k+1}\rho_{2k+1}}$. A little calculation shows that this can be written as
$\bar{\Theta}_{\mu\nu}=-\frac{1}{(2k)!}G_{\mu}\bar{\cdot}G_{\nu}+\frac{1}{4}\frac{1}{(2k)!}(dP-\bar{\star}dP)_{\mu}\bar{\cdot}(dP-\bar{\star}dP)_{\nu}+\bar{\Theta}^{\prime}_{\mu\nu}.$ (146)
@@
To fix this we need to consider $\Upsilon\neq 0$ and include the last term in (A). This gives an additional variation
$\delta^{\prime}S$ $=\int\frac{1}{2}\Upsilon\wedge(dP-\bar{\star}dP)+\frac{1}{2}\Upsilon\wedge M(Q)-\frac{1}{8}(\bar{\delta}\bar{\star}){M}(Q)\wedge M(Q).$ (147)
A natural choice is $\Upsilon=\tfrac{1}{2}(\bar{\delta}\bar{\star})M(Q)$, which we have seen is self-dual, leading to
@@
$=(\bar{\delta}\bar{\star})F,$ (148)
and
$\delta^{\prime}S$ $=-\int\frac{1}{4}(dP-\bar{\star}dP)\wedge(\bar{\delta}\bar{\star})M(Q)+\frac{1}{8}M(Q)\wedge(\bar{\delta}\bar{\star})M(Q).$ (149)
@@
$\bar{\Theta}^{\prime}_{\mu\nu}=-\frac{1}{4}\frac{1}{(2k)!}\left((dP-\bar{\star}dP)_{\mu}\bar{\cdot}M(Q)_{\nu}+(dP-\bar{\star}dP)_{\nu}\bar{\cdot}M(Q)_{\mu}+M_{\mu}(Q)\bar{\cdot}M_{\nu}(Q)\right),$ (150)
@@
Here we see that the first term is conserved on-shell, i.e. it’s divergence with respect to $\bar{\nabla}$ vanishes on-shell, whereas the second term vanishes on-shell. As a result, $\bar{\nabla}^{\mu}\bar{\Theta}_{\mu\nu}=0$ on-shell.
@@
The action is invariant under such diffeomorphisms, up to possible boundary terms which we will discard. We observe that